Compute the value of the discriminant and give the number of real solutions to the quadratic equation.
-4x^2 + 4x - 1 = 0

Answer by stanbon(57219)

Compute the value of the discriminant and give the number of real solutions to the quadratic equation.
-4x^2 + 4x - 1 = 0
----
Discriminant = 4^2-4*-4*-1 = 16 -16 = 0
-----
Two Real equal solutions.
Cheers,
Stan H.